How many atoms are in a 250 g sample of Gold?

A. 1.19 × 1023 atoms
B. 1.19 × 1026 atoms
C. 7.65 × 1022 atoms
D. 7.65 × 1024 atoms
E. None of the Above
1
Expert's answer
2018-01-18T13:49:07-0500
m(Au)=250 g
n(Au)=m/M=250g/197g/mol = 1,27 moles
N=n•Na
N(Au) = 1,27•6,02•1023 = 7,65•1023 atoms
E. None of the Above